# Revisiting Stochastic Gravitational-wave Background in the Strong Signal Case

### v1

@@ VI Conclusions and discussions
In this paper, we have updated the SNR tailored for detecting loud SGWB. By utilizing the improved SNR estimation, we have plotted PLIS and PIS curves for TianQin, the TianQin I+II network, and TianQin + LISA network to assess the detectable SGWB energy spectrum density over an operation time of 1 year. Our findings indicated that for SNR values below 10, the impact of considering or neglecting the weak-signal limit on the results is minimal. However, when the SNR exceeds 100, applying the weak-signal limit can lead to a considerable overestimation of the detectable energy spectrum density. This overestimation becomes markedly pronounced for the TianQin + LISA network, where it may be inflated by approximately an order of magnitude.
Additionally, we have employed SNR contour plots to determine the detectable parameter regions for loud SGWBs generated by EDWDs and first-order PTs. Our analysis revealed a significant discrepancy, especially concerning the TianQin + LISA network, when comparing results obtained with and without the weak-signal limit using a selected SNR threshold of 100. Under the weak-signal limit, the potential for detecting the SGWB from EDWDs of DA and DB types appears promising within a one-year operation time, maintaining a consistent merger rate of about $500\,\,{\rm Mpc^{-3}\,Myr^{-1}}$ across different redshifts. However, this merger rate is notably below the required threshold as the extragalactic SGWB is strong enough. Similarly, misestimations arise in constraining the parameters of first-order PTs, where uncertainties in the PT models only allow for broad constraints on the detectable parameter regions.
We emphasize that our analysis can extend to the statistical inference of SGWB detection, which encompasses classical (frequentist) inference and Bayesian inference. The pivotal aspect lies in constructing the likelihood function. By estimating the expectation and variance in terms of Eq. (24) and Eq. (25), one can ensure that the likelihood function aligns more closely with the actual detection process, thereby minimizing bias in parameter estimation. This approach is not limited to space-borne detectors Adams and Cornish (2014); Caprini et al. (2019); Pieroni and Barausse (2020); Flauger et al. (2021), but is also relevant for next-generation ground-based detectors Sachdev et al. (2020); Zhou et al. (2023); Zhong et al. (2023); Pan and Yang (2023); Bellie et al. (2023); Song et al. (2024). Ongoing work on statistical inference regarding this matter is underway and will be presented elsewhere.
@@ Acknowledgements.
This work has been supported by the Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2019B030302001), the National Key Research and Development Program of China (No. 2020YFC2201400), the Natural Science Foundation of China (Grants No. 12173104), the Natural Science Foundation of Guangdong Province of China (Grant No. 2022A1515011862), and the Guangdong Basic and Applied Basic Research Foundation(Grant No. 2023A15150301). Z.C.L. is supported by the China Postdoctoral Science Foundation (Grant No. 2023M744094). We also thank Fapeng Huang, Yun Jiang, and Jianwei Mei for helpful discussions.

### v2

@@ VI Conclusions and discussions
In this paper, we have derived the SNR tailored for detecting loud SGWB. By utilizing the improved SNR estimation, we have plotted PLIS and PIS curves for TianQin, the TianQin I+II network, and TianQin + LISA network to assess the detectable SGWB energy spectrum density over an operational period of 1 year. Our findings indicated that for SNR values below 10, the impact of considering or neglecting the weak-signal limit on the results is minimal. However, when the SNR exceeds 100, applying the weak-signal limit can lead to a considerable overestimation of the detectable energy spectrum density. This overestimation becomes markedly pronounced for the TianQin + LISA network, where it may be inflated by approximately an order of magnitude.
Additionally, we have employed SNR contour plots to determine the detectable parameter regions for loud SGWBs generated by EDWDs and first-order PTs. With a chosen SNR threshold of 100, the prospect for detecting the SGWB from EDWDs, specifically DA and DB types, appears promising. This detection is grounded in a consistent merger rate exceeding $200\,\,{\rm Mpc^{-3}\,Myr^{-1}}$ surpassing varying redshifts, with the operation period extending over 1 year. On the other hand, while it is feasible to narrow down the parameters for first-order PTs by holding certain variables fixed, uncertainties in the underlying models permit only a broad constriction of the detectable parameter region.
We emphasize that our analysis can extend to the statistical inference of SGWB detection, which encompasses classical (frequentist) inference and Bayesian inference. The pivotal aspect lies in constructing the likelihood function. By estimating the expectation and variance in terms of Eq. (25) and Eq. (26), one can ensure that the likelihood function aligns more closely with the actual detection process, thereby minimizing bias in parameter estimation. This approach is not limited to space-borne detection Caprini et al. (2019); Pieroni and Barausse (2020); Flauger et al. (2021), but is also relevant for next-generation ground-based detection Zhong et al. (2023); Bellie et al. (2023).
In this work, we assume the ideal case where the only GW signal is SGWB. In reality, different signals will overlap and make the analysis more complicated. Fortunately, even if the SGWB is strong, it will not complicate the analysis of other sources. Taking the foreground from Galactic compact binaries as an example, it is expected that this foreground is comparable to or even stronger than instrument noise. However, when analyzing signals like the merger of massive black holes, the foreground can just be regarded as part of the PSD. That being said, the strong SGWB will inevitable raise the noise level, causing a decrease in the sensitivity.
@@ Acknowledgements.
This work has been supported by the Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2019B030302001), the National Key Research and Development Program of China (No. 2020YFC2201400), the Natural Science Foundation of China (Grants No. 12173104), the Natural Science Foundation of Guangdong Province of China (Grant No. 2022A1515011862), and the Guangdong Basic and Applied Basic Research Foundation(Grant No. 2023A1515030116). Z.C.L. is supported by the China Postdoctoral Science Foundation (Grant No. 2023M744094), and the Guangdong Basic and Applied Basic Research Foundation(Grant No. 2023A1515111184). We also thank Fapeng Huang, Yun Jiang, and Jianwei Mei for helpful discussions.